OpenStudy (anonymous):

how would you graph y+x=2

OpenStudy (anonymous):

first find x when y=0 then y when x=0 so this gives you the x and y intercepts

OpenStudy (anonymous):

so when y=0 in y+x=2 0+x=2 therefore x=2

OpenStudy (anonymous):

also when x=0 in y+x=2 y+0=2 y=2 so you have the coordinates (2,2) now graph

OpenStudy (anonymous):

k thank you

OpenStudy (anonymous):



OpenStudy (anonymous):

but you need to draw to scale to have better accurate graph

OpenStudy (anonymous):

k